5x-7y=-21
14y-5x=22
using substitution or elimination

Answer by MathTutorer(13)

5x-7y=-21
14y-5x=22
Get 5x by its self
5x = 7y - 21 (after adding 7y to both sides)
Factor that into the second equation
14y - (7y - 21) = 22
Then simplify
7y - 21 = 22
then simplify again
7y = 43
y = 43/7